curriful:

What's the new Y-Intercept of 1=-3/5x+b

NaiNoah:

(0, - 5/3)

toga:

To find the y-intercept of the equation 1=-3/5x+b, we need to isolate the constant term 'b'. First, we can move the variable term '-3/5x' to the right-hand side of the equation by adding it to both sides. This gives us: 1 + 3/5x = b Now we have the equation in the form y = mx + b, where 'm' is the slope and 'b' is the y-intercept. So, we can see that the y-intercept of the given equation is b = 1 + 3/5x.

curriful:

@toga wrote:
To find the y-intercept of the equation 1=-3/5x+b, we need to isolate the constant term 'b'. First, we can move the variable term '-3/5x' to the right-hand side of the equation by adding it to both sides. This gives us: 1 + 3/5x = b Now we have the equation in the form y = mx + b, where 'm' is the slope and 'b' is the y-intercept. So, we can see that the y-intercept of the given equation is b = 1 + 3/5x.
thanks!

toga:

@curriful wrote:
@toga wrote:
To find the y-intercept of the equation 1=-3/5x+b, we need to isolate the constant term 'b'. First, we can move the variable term '-3/5x' to the right-hand side of the equation by adding it to both sides. This gives us: 1 + 3/5x = b Now we have the equation in the form y = mx + b, where 'm' is the slope and 'b' is the y-intercept. So, we can see that the y-intercept of the given equation is b = 1 + 3/5x.
thanks!
np